OpenStudy (anonymous):

A certain kind of candy comes in mixed bags of five shapes: 40% are flowers, 30% are snowflakes, 20% are butterflies, and 10% are diamonds. Luwanda's favorites are the flowers and the butterflies. What is the probability that the next one she picks (at random) will be one of her favorites?

OpenStudy (texaschic101):

flowers = 40/100 which reduces to 2/5 butterflies = 20/100 which reduces to 1/5 2/5 * 1/5 = 2/25 probability you might better get a second opinion because I am not the best at these

OpenStudy (anonymous):

me either :( but thank you